OpenStudy (anonymous):

1/3 + 1/9 + 1/27 + 1/81 + 1/243 + ...

OpenStudy (tkhunny):

This kind of series has a name and known properties. Identification is the first, most important part. What kind of series is it?

OpenStudy (anonymous):

it looks like a geometric sequence

OpenStudy (tkhunny):

That was my impression. Normally: 'a' is the first term and 'r' is the common ratio. Can you identify those parts?

OpenStudy (anonymous):

a=1/3 and r = 1/3

OpenStudy (tkhunny):

That's it. Now what's the sum of such an infinite series?